If 3 times Jane's age, in years, is equal to 8 times Beth's age, in years, and the difference between their ages is 15 years, how old are Jane and Beth?
Answer by edjones(8007)

Let j=jane's age and b=beth's age
3j=8b
j-b=15
.
j=b+15
3(b+15)=8b
3b+45=8b
5b=45
b=9
j=24
.
Ed